dmitri plays hockey. He earns 1 point for every goal he scores and 1 point for every assist. This season he earned 63 points. He scored 17 fewer goals than assist. How many goals did dmitri score this season? How many assist did he have?
Answer by Theo(13342)

g = number of goals
a = number of assists.
g + a = 63
g = a - 17
replace g with a - 17 to get:
a - 17 + a = 63
combine like terms to get:
2a - 17 = 63
add 17 to both sides to get:
2a = 80
solve for a to get:
a = 80/2 = 40
g = 40 - 17 = 23
g + a = 23 + 40 = 63
number of goals is 17 fewer that number of assists.
number of goals plus number of assists is equal to 63.
solution is he scored 40 assists.